(9 pts) Suppose the random variable has a geometric distribution with parameterp. Let V = Find the probability distribution of V. 4.
Answer #1

The cumulative distribution of V is

F_V(v) = P(Vleq v) = P(rac{Y^{rac{1}{3}}}{3} leq v)

Rightarrow F_V(v) = P(rac{Y^{rac{1}{3}}}{3} leq v)

Rightarrow F_V(v) = P(Y^{rac{1}{3}} leq 3v)

Rightarrow F_V(v) = P(Y leq 27v^3)

Rightarrow F_V(v) = F_y(27v^3)

The probability density function of V is

f_V(v) = rac{partial }{partial v}F_V(v) = rac{partial }{partial v}F_y(27v^3) = 27*3*v^2*f_y(27v^3)

fy(1) 27 * 3 * 1,2 * fy(271,3)

Rightarrow f_V(v) =81v^2*p*(1-p)^{27v^3 - 1}